OpenStudy (anonymous):

2, 5, 10, 17, . . . Which equation represents the rule for finding the nth term of this sequence?

OpenStudy (welshfella):

Hint: it includes n^2

OpenStudy (welshfella):

lets square the numbers n = 1,2,3... N = 1 = 1 n= = 2^2 = 4 n = 3 = 3^2 = 9 n=4 4^2 = 16 so we have 1, 4, 9, 16 compare this with the sequence in the question Do you see the pattern?